Solve the inequality algebraically.
7x-6>=-3x^2
The solution set is __?
(Type your answer in interval notation. Use integers or fractions for any numbers in the expression. Simplify your answer.)

Answer by blacksaibot(4)

3x^2 + 7x - 6 >= 0
(3x-2)(x+3)

3x-2 = 0
x = (2/3)

x+3 = 0
x = -3

Solutions: x = (2/3) and -3